# Orchestrated Robust Controller for the Precision Control of Heavy-duty Hydraulic Manipulators

### v1

@@ VII-C Experimental Result
Finally, experimental results are juxtaposed to those reported in [5], utilizing $\rho$ value as a metric for comparative assessment. Table IV exhibits the computed performance index value for works in which the high-DoF model of HHM is considered. As can be seen, the proposed method has achieved a better $\rho$ value in comparison to other works, demonstrating its excellent behavior in Cartesian space motion control.
Considering all the experimental results provided in this section, along with the comparison to the original VDC approach, the validity of the theoretical results is verified. As shown, tackling unknown model and actuator uncertainties with RBFNNs and handling compound input constraints with an adaptive deadzone-backlash controller significantly improved the tracking performance, both in transient and steady-state. As it was mathematically proved in Theorem 3 and experimentally shown in this section, the designed controller ensured SGUUB stability for the system and converged the tracking error to a substantially small neighborhood of origin.
@@
Study
$\rho$
DoF
$\rho$/DoF
This study
0.015 6 0.0025
Kalmari 2015
0.12
4
0.03
Nguyen 2010
0.32
4
0.08
Zhu 2005
0.015
5
0.003
Tsukamoto 2002
0.126
6
0.021
Egelan 1987
0.038
8
0.0047
@@ VIII Conclusion
In this study, an orchestrated robust controller is designed to achieve a precise and robust performance. For this purpose, VDC was employed as the baseline controller. VDC decomposes the entire complex system into subsystems of the rigid body and actuators by means of VCP. Then, the controller design and stability analysis are accomplished at the subsystem level and expanded to the entire system by means of VPFs and virtual stability. A novel way of incorporating RBFNNs and adaptive inverse deadzone-backlash controllers into each subsystem resulted in the orchestrated robust controller approach. Considering all unknown model and actuator uncertainties, compound input constraints, and unknown disturbances, the SGUUB stability is achieved for the first time in the context of VDC. All theoretical results are validated via extensive simulation and experimental results concerning the highly nonlinear, six DoF heavy-duty hydraulic manipulator. The simulation results demonstrated that the proposed controller outperformed the original VDC, PD, and the state-of-the-art method in the presence of compound input constraints. Moreover, the experimental results demonstrate that the proposed method tackles unknown model and actuator uncertainties and reduces the impact of compound input constraints in real-world applications, resulting in considerably lower transient and steady-state tracking errors and lower actuator force errors in comparison to the original VDC for a given power. Additionally, the proposed method ended up with a precise tracking performance in Cartesian space in comparison to the original VDC. All the analyses, including sensitivity to control gain, tracking performance in joint space and Cartesian space, comparison to the state-of-the-art controller, performance evaluation in the sense of $\rho$ value, and implementation on the real-world six DoF commercial manipulator, display the universality of the proposed method in practical applications.

### v2

@@ VII-C Experimental Result
Finally, experimental results are juxtaposed to those reported in [5], utilizing $\rho$ value as a metric for comparative assessment. Table IV exhibits the computed performance index value for all the works along with its normalization based on the end-effector DoF (EE-DoF), indicating number of DoFs controlled in Cartesian space with 3 DoF for position and 3-DoF for orientation. As can be seen, the proposed method has achieved a better $\rho$ value in comparison to other works with full pose control, demonstrating its excellent behavior in Cartesian space motion control.
Considering all the experimental results provided in this section, along with the comparison to the original VDC approach, the validity of the theoretical results is verified. As shown, tackling unknown model and actuator uncertainties with RBFNNs and handling compound input nonlinearities with an adaptive deadzone-backlash controller significantly improved the tracking performance, both in transient and steady-state. As it was mathematically proved in Theorem 3 and experimentally shown in this section, the designed controller ensured SGUUB stability for the system and converged the tracking error to a substantially small neighborhood of origin.
@@
Study $\rho$ EE-DoF $\rho$/EE-DoF
This study 0.015 6 0.0025
Koivumaki 2019 [14] 0.0057 2 0.0029
Kalmari 2015 0.12 3 0.04
Zhu 2005 0.015 3 0.005
Tsukamoto 2002 0.126 6 0.021
Egelan 1987 0.038 3 0.012
@@ VIII Conclusion
In this study, an orchestrated robust controller is designed to achieve a precise and robust performance. For this purpose, VDC was employed as the baseline controller. VDC decomposes the entire complex system into subsystems of the rigid body and actuators by means of VCP. Then, the controller design and stability analysis are accomplished at the subsystem level and expanded to the entire system by means of VPFs and virtual stability. A novel way of incorporating RBFNNs and adaptive inverse deadzone-backlash controllers into each subsystem resulted in the orchestrated robust controller approach. Considering all unknown model and actuator uncertainties, compound input nonlinearities, and unknown disturbances, the SGUUB stability is achieved for the first time in the context of VDC. All theoretical results are validated via extensive simulation and experimental results concerning the highly nonlinear, six DoF HHM. The simulation results demonstrated that the proposed controller outperformed the original VDC, PD, and the state-of-the-art method in the presence of compound input nonlinearities. Moreover, the experimental results demonstrate that the proposed method tackles unknown model and actuator uncertainties and reduces the impact of compound input nonlinearities in real-world applications, resulting in considerably lower transient and steady-state tracking errors and lower actuator force errors in comparison to the original VDC for a given power. Additionally, the proposed method ended up with a precise tracking performance in Cartesian space in comparison to the original VDC. All the analyses, including sensitivity to control gain, tracking performance in joint space and Cartesian space, comparison to the state-of-the-art controller, performance evaluation in the sense of $\rho$ value, and implementation on the real-world generic manipulator, display the universality of the proposed method in theory and applications.